OpenStudy (anonymous):

Find dy/dx when x = 1 , given that y = (x^2 + 3) ( 2 - 4x - 5x^2)

OpenStudy (anonymous):

product rule = f(x)g'(x) + g(x)f'(x) (x^2+3)(-10x -4) + (-5x^2 -4x +2)(2x) = (-10x^3 - 4x^2 - 30x -12) + (-10x^3 - 8x^2 +4x) = -20x^3 -12x^2 -26x -12 Take out a (-2) -2(10x^3 + 6x^2 + 13x + 6) add in x = 1 -2( 10 + 6 + 13 + 6) = -2(35) (dy/dx) = -70 when x = 1

OpenStudy (anonymous):

wow thank you so much

OpenStudy (anonymous):

no problem